OpenStudy (anonymous):

A filter circuit is activated when the current reaches 5 amperes. If a charge is fed to the circuit starting at 1 ampere and is steadily increased by 50% of the current value every 10 seconds, how long will it be until the filter activates? 3.69 seconds 10.50 seconds 39.69 seconds 40.00 seconds

OpenStudy (yttrium):

does it something to do with integration?

ganeshie8 (ganeshie8):

final value = 5 amperes initial value = 1 ampere rate = 1 + 0.5 = 1.5 every 10 seconds set up a growth equaiton

ganeshie8 (ganeshie8):

\(\large 5 = (1.5)^{\frac{t}{10}}\)

ganeshie8 (ganeshie8):

solve \(t\)